1.Calculate the pH and pOH a solution if the concentration of hydrogen ions is equal to 5.35103 mole/L.

1
Expert's answer
2022-02-14T08:15:04-0500

"pH = - log [H3O+]."

pH= - log(5.35× 103)

pH= 3.7

pOH= 14-3.7= 10.3